OpenStudy (anonymous):

please help with SERIES problem.. \[\sum_{n=1}^{\infty} \frac{ (-1)^\left( n+1 \right) }{ 2^n+n }\] (a) estimate the series by evaluating \[S_2\] i put n=1,2 so S_2 \[\frac{ 1 }{ 3 }-\frac{ 1 }{ 5 }=\frac{ 2 }{ 15 }=S_2\] (is this correct?) (b) determine the error bound for \[\left| S_3-\sum_{n=1}^{\infty} \frac{ (-1)^\left( n \right) }{ 2^n+n } \right|\]

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ (-1)^\left( n+1 \right) }{ 2^n+n }\] (a) estimate the series by evaluating \[S_2\] i put n=1,2 so S_2 \[\frac{ 1 }{ 3 }-\frac{ 1 }{ 5 }=\frac{ 2 }{ 15 }=S_2\] (is this correct?) (b) determine the error bound for \[\left| S_3-\sum_{n=1}^{\infty} \frac{ (-1)^\left( n \right) }{ 2^n+n } \right|\]

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ (-1)^\left( n+1 \right) }{ 2^n+n }\] (a) estimate the series by evaluating \[S_2\] i put n=1,2 so S_2 \[\frac{ 1 }{ 3 }-\frac{ 1 }{ 5 }=\frac{ 2 }{ 15 }=S_2\] (is this correct?) (b) determine the error bound for \[\left| S_3-\sum_{n=1}^{\infty} \frac{ (-1)^\left( n \right) }{ 2^n+n } \right|\]

OpenStudy (anonymous):

(b) part please..

OpenStudy (loser66):

you mean we have to figure out the error from that question?

OpenStudy (anonymous):

yes it's all related problem

OpenStudy (anonymous):

is (a) part correct?

OpenStudy (amistre64):

if i recall correctly, the error is bound between the average of\[\int_{n}^{\infty}f~and\int_{n+1}^{\infty}f\]

OpenStudy (anonymous):

could you please to go over step by step.?

OpenStudy (loser66):

@amistre64 don't understand, please, more explanation

OpenStudy (amistre64):

the notation goes something like this \[S_\infty=S_n+S_{n+1}\] the whole is the sum of the parts; if you can sum up to a finite term, the remainder can be approximated as the average of n to inf, and n+1 to inf

OpenStudy (amistre64):

the bounds tho ... thats fuzzy

OpenStudy (amistre64):

this might do a better job at explaining the detials :) http://tutorial.math.lamar.edu/Classes/CalcII/EstimatingSeries.aspx

OpenStudy (anonymous):

made me more confused.. haha

OpenStudy (anonymous):

how do i approach first? to determine the error bound?

OpenStudy (amistre64):

my picture seems to have gotten lost ....

OpenStudy (amistre64):

essentally; the area under f from n+1 to infinity is the lower bound and the area under f from n to infinity is the upper bound\[\int_{n+1}^{inf}f<R_n<\int_{n}^{inf}f\]

OpenStudy (amistre64):

you determine them by taking the limit as n to infinity of the integral of f

OpenStudy (amistre64):

the alternating part seems a bit awkward to me, but thats the principle to what ive learned

OpenStudy (anonymous):

ok what's the next step...??

OpenStudy (amistre64):

examples 2 and 3 of the link i posted seem like they are most apt to fit this problem

OpenStudy (amistre64):

\[\left|\frac{(-1)^\left( n+1 \right) }{ 2^n+n }\right|< (\frac{ 1 }{ 2 })^n\]